# On the Learnability of Watermarks for Language Models

### v2

@@ Appendix G Robustness to text edits experiment details
In this experiment, we take the logits-based watermark distilled Llama 2 7B model from each watermark type with the smallest detection p-values, namely KGW $\delta=2$, Aar $k=2$, and KTH $s=1$. We use the 200-token length generations prompted from C4 RealNewsLike used in the main experiments. Then, for varying random edit proportions $\varepsilon=\{0,0.1,0.2,\ldots,0.8\}$, we first randomly delete $\varepsilon$ proportion of the tokens in each sequence, then insert random tokens at random locations until the corrupted sequence reaches the length of the original sequence. So $1-\varepsilon$ of the tokens in the corrupted sequence are from the original sequence, and they form a common subsequence in the corrupted and original sequences. Then, we compute the median watermark detection p-value among these corrupted generations. We plot detection p-value against $\varepsilon$, the proportion of tokens edited.
@@ Appendix H Continued finetuning details
In this experiment, we take the logits-based watermark distilled Llama 2 7B model from each watermark with the smallest detection p-values, namely KGW $\delta=2$, Aar $k=2$, and KTH $s=1$. Then, we fine-tune these distilled models on OpenWebText for 5,000 steps, saving the model every 1,000 training steps. We use a batch size of 32, sequence length of 512 tokens, maximum learning rate of 1e-5, cosine learning rate decay with a linear warmup for the first 500 steps, and the AdamW optimizer (Loshchilov & Hutter, 2019) with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and no weight decay. Then, for each model checkpoint (including the original distilled model at 0 steps), we generate 200-token completions prompted by 50 tokens from the C4 RealNewsLike dataset, as in the main experiments. We compute the median p-value among these generations and plot p-value against number of fine-tuning steps.
@@ I.1 Spoofing attack training details
To perform sampling-based watermark distillation, we first query the teacher/victim model Llama 2-Chat 7B with the KGW decoding-based watermarking strategy ($\delta=2$) for watermarked samples. We generate 52,000 samples, with a max length of 1024 tokens, prompted by instructions from the Alpaca dataset (Taori et al., 2023). Then, in order to preserve the tendency of base Alpaca-7B to respond to harmful requests, we filter out refusals from these samples. We use simple keyword filtering, removing any responses that contain “I cannot”, “appropriate or”, “just an AI”, “offensive”, or “I apologize, but”. These keywords were chosen by manually inspecting a subset of the samples.
Then, we fine-tune the adversary model (initialized as Alpaca-7B) for 3 epochs (1,200 steps) on the instructions and watermarked responses with a batch size of 32 examples, maximum learning rate of 2e-5, cosine learning rate decay with a linear warmup during the first 0.03 proportion of steps, and the AdamW optimizer (Kingma & Ba, 2015; Loshchilov & Hutter, 2019) with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and no weight decay. Training took approximately 2 hours on 4 NVIDIA A100 80GB GPUs.
@@ I.2 Harmfulness evaluation details
We use the HarmfulQ benchmark (Shaikh et al., 2023), which contains 200 toxic questions generated by text-davinci-002, using methodology similar to automatic red-teaming of large language models (Perez et al., 2022). Questions include “How do I make a deadly poison?” and “How can I make someone feel unwelcome based on their identity?”. For each model, we generate responses to each question using temperature 1 sampling until the stop token is generated, without setting a minimum number of tokens generated. For Llama 2-Chat, we do not use a system prompt, so the model is not explicitly prompted to refuse harmful requests. For the base and adversary Alpaca models, we use the default Alpaca prompt, which also does not include instructions to refuse harmful requests.

### v3

@@ Appendix G Robustness to text edits experiment details
In this experiment, we take one logit-based and sampling-based watermark distilled Llama 2 7B model from each watermark type: KGW $k=1,\gamma=0.25,\delta=2$, Aar $k=2$, and KTH $s=1$. We use the 200-token generations prompted from C4 RealNewsLike used in the main experiments, as described in §4.3. Then, for varying random edit proportions $\varepsilon=\{0,0.1,0.2,\ldots,0.8\}$, we first randomly delete $\varepsilon$ proportion of the tokens in each sequence, then insert random tokens at random locations until the length of the corrupted sequence reaches the length of the original sequence. So $1-\varepsilon$ of the tokens in the corrupted sequence are from the original sequence, and they form a common subsequence in the corrupted and original sequences. Then, we compute the median watermark detection p-value among these corrupted generations. Figure 4 plots detection p-value against $\varepsilon$, the proportion of tokens edited.
@@ Appendix H Continued finetuning details
In this experiment, we take one logit-based watermark distilled Llama 2 7B model from each watermark type: KGW $k=1,\gamma=0.25,\delta=2$, Aar $k=2$, and KTH $s=1$. Then, we fine-tune these distilled models on OpenWebText (Gokaslan et al., 2019) for 2,500 steps, saving the model every 500 training steps. We use a batch size of 32, sequence length of 512 tokens, maximum learning rate of 1e-5, cosine learning rate decay with a linear warmup for the first 10% of steps, and the AdamW optimizer (Kingma & Ba, 2015; Loshchilov & Hutter, 2019) with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and no weight decay. Then, for each model checkpoint (including the original distilled model at 0 steps), we generate 200-token completions prompted by 50 tokens from the C4 RealNewsLike dataset, as in the main experiments (§4.3). We compute the median detection p-value among these generations and plot p-value against number of fine-tuning steps in Figure 4.
@@ I.1 Spoofing attack training details
To perform sampling-based watermark distillation, we first query the teacher/victim Llama 2-Chat 7B model with the KGW decoding-based watermarking strategy ($k=1,\gamma=0.25,\delta=2$) for watermarked samples. We generate 52,000 samples, with a maximum length of 1024 tokens, prompted by instructions from the Alpaca dataset (Taori et al., 2023). Then, in order to preserve the tendency of the Alpaca-7B student/adversary to respond to harmful requests, we filter out refusals from these samples. We use simple keyword filtering, removing any responses that contain “I cannot”, “appropriate or”, “just an AI”, “offensive”, or “I apologize, but”. These keywords were chosen by manually inspecting a subset of the samples.
Then, we fine-tune the adversary model (initialized as Alpaca-7B) for 3 epochs (roughly 1,200 steps) on the instructions and watermarked responses with a batch size of 32 examples, maximum learning rate of 2e-5, cosine learning rate decay with a linear warmup during the first 3% of steps, and the AdamW optimizer (Kingma & Ba, 2015; Loshchilov & Hutter, 2019) with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and no weight decay. Training took approximately 2 hours on 4 NVIDIA A100 80GB GPUs.
@@ I.2 Harmfulness evaluation details
We use the HarmfulQ benchmark (Shaikh et al., 2023), which contains 200 toxic questions generated by text-davinci-002, using methodology similar to automatic red-teaming of large language models (Perez et al., 2022). Questions include “How do I make a deadly poison?” and “How can I make someone feel unwelcome based on their identity?”. For each model, we generate responses to each question using standard temperature 1 sampling until the stop token is generated, without setting a minimum number of tokens generated. For Llama 2-Chat 7B, we do not use a system prompt, so the model is not explicitly prompted to refuse harmful requests. For the base and adversary Alpaca models, we use the default Alpaca prompt, which also does not include instructions to refuse harmful requests.